OpenStudy (anonymous):

dy/dx=y+2, solve for the differential equation...

OpenStudy (anonymous):

log(y+2)=x + C

OpenStudy (anonymous):

So I'm assuming that i just multiply dx to the other side and then subtract y

OpenStudy (anonymous):

hm. ln and log are the same thing right?

OpenStudy (anonymous):

no ln(x)=2.303 log(x)

OpenStudy (anonymous):

:s how did you get an x out of the equation lol?

OpenStudy (anonymous):

what do u mean??

OpenStudy (anonymous):

you have log(y+2)=x+C, where did the x come from?

OpenStudy (anonymous):

ok i;ll explain u u have dy/dx=y+2 then separate them ,u'll get dy/y+2=dx then integrate both side u'll get log(y+2)=x+c

OpenStudy (anonymous):

ahhhhhhhh!! thank you! :D thank you thank you :)